### v1

@@ 5.1.2 Numerical performance of Perishing-Guardrail
Figure 5: Algorithm comparison across $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right],\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right],\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$, and $\mathbb{E}\mathopen{}\mathclose{{}\left[\text{Spoilage}}\right]$, for $\alpha\in\{0.1,0.2,0.25,0.3\}$.
$\alpha$ Static-$\frac{B}{\overline{N}}$ Static-$\underline{X}$ Vanilla-Guardrail Perishing-Guardrail
.1.1.1.1 $0.99\pm 0.020$ $0.00\pm 0$ 0.0 $1.00\pm 0.0$ $0.11\pm 0.06$
.2.2.2.2 $0.63\pm 0.095$ $0.00\pm 0$ 0.0 $0.68\pm 0.091$ $0.03\pm 0.03$
.3.3.3.3 $0.03\pm 0.037$ $0.00\pm 0$ 0.0 $0.06\pm 0.046$ $0.00\pm 0$ 0.0
@@
  • High Perishing ($\alpha=0.1$): While unfairness (as measured by $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$ and $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$) is decreasing in $T$ under Perishing-Guardrail and Static-$\underline{X}$, Vanilla-Guardrail and Static-$\frac{B}{\overline{N}}$ perform remarkably poorly along these two metrics. This is due to the fact that these latter algorithms fail to account for the unavoidable perishing loss, resulting in an extremely high stockout probability, as illustrated in Table 1, for $T=150$. In contrast, the two perishing-aware algorithms rarely run out of resources. This underscores the importance of modifying the baseline guardrail $\underline{X}$, which was specifically constructed to avoid stockouts due to unavoidable perishing.
    Comparing Static-$\underline{X}$ to Perishing-Guardrail, our results also demonstrate that, in this high-perishing regime, the strategy of cautiously over-allocating by $L_{T}$ comes at a significant reduction in inefficiency $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right]$, at close to no increase in counterfactual envy $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$.
  • Medium Perishing ($\alpha=0.2$): Though we observe similar trends as when $\alpha=0.1$, all algorithms perform better across the board. Still, in Table 1 we see that the perishing-agnostic algorithms run out of resources in over 50% of replications. As observed in Fig. 3, Perishing-Guardrail exhibits both higher efficiency and lower spoilage than its perishing-agnostic counterpart in this regime, since it satisfies the threshold condition more frequently, as described above.
  • Low Perishing ($\alpha=0.3$): For this smaller perishing rate, Vanilla-Guardrail stocks out significantly less frequently. Putting this together with the fact that $B/\overline{N}>\underline{X}$, this explains the fact that it has lower counterfactual envy than Perishing-Guardrail. However, along all other metrics Perishing-Guardrail improves upon Vanilla-Guardrail. The improvements in efficiency and spoilage are due to the same effects as described above; moreover, our algorithm improves upon Vanilla-Guardrail on $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$ since it never stocks out.
Overall, our results highlight the robustness of Perishing-Guardrail to perishability, as it is able to achieve similar if not improved performance as Vanilla-Guardrail in settings where there is limited perishing, with vastly superior performance in high-perishing settings.
Algorithm $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Spoilage}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Stockout}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right]$
@@
Table 2: Performance of the different algorithms (for $L_{T}=T^{-0.35}$) on the “ginger” dataset in Keskin et al., (2022). The second number in each cell corresponds to 95% confidence intervals.
Real-World Instance.
We next investigate the performance of our algorithm using the “ginger” dataset provided by Keskin et al., (2022), which tracks demand, replenishments, and perishing of ginger across $T=365$ days. We treat the time between each replenishment as an independent sample (102 samples in total), and fit a geometric distribution to the quantity of goods that perish in each sample, obtaining $p=0.00224$. We similarly fit a truncated normal distribution to the dataset, with $\mathcal{N}(3.2,1.85)$. Finally, we let $B=365\cdot 3.2=1168$. For these inputs, $B/\overline{N}=0.89$ and $\underline{X}=0.46$. Under these parameters the offset-expiry condition is only satisfied 65.2% of the time; given this aggressive perishing, perishing-agnostic algorithms are expected to perform particularly poorly.
@@
  • As conjectured, Static-$\frac{B}{\overline{N}}$ and Vanilla-Guardrail stock out on 100% of replications since they fail to account for endogenous perishing. The high stockout probabilities of Static-$\frac{B}{\overline{N}}$ and Vanilla-Guardrail lead to high unfairness (vis-à-vis hindsight and counterfactual envy), since later arrivals receive allocations of zero. In contrast, Static-$\underline{X}$ never stocks out. Perishing-Guardrail achieves a higher stockout rate of 40%, likely due to the fact that threshold condition does not account for non-offset-expiring trajectories. Still, our algorithm’s counterfactual envy and hindsight envy are over 30% and 70% lower, respectively, than that of Vanilla-Guardrail.
  • Perishing-Guardrail allocates approximately 10% fewer goods than Vanilla-Guardrail. It is notable, however, that it is more efficient than Static-$\frac{B}{\overline{N}}$; this highlights that naively allocating more aggressively need not always generate gains.
Overall, we see that even when offset-expiry holds with much lower probability, for small losses in efficiency our algorithm makes major gains in fairness relative to perishing-agnostic algorithms.
@@
As seen in Section 3, the performance of any algorithm is a function of $\mathcal{L}^{\textsf{perish}}$, which depends on the allocation order $\sigma$ in non-i.i.d. settings. A natural question that arises, then, is how a decision-maker should choose $\sigma$. In this section, we investigate the impact of three practical allocation orders on the numerical performance of Perishing-Guardrail.
Given Theorem 3.7, a reasonable allocation order to choose would be $\sigma^{*}\in\arg\min_{\sigma}\mathcal{L}^{\textsf{perish}}(\sigma),$ where we emphasize the dependence of $\mathcal{L}^{\textsf{perish}}$ on the order $\sigma$. Computing such a $\sigma^{*}$, however, is infeasible given the space of $N!$ orderings. In lieu of this, we identify sufficient conditions on the perishing process and allocation order that guarantee that $\mathcal{L}^{\textsf{perish}}$ remain low (equivalently, that $\underline{X}$ remain high), and use these conditions to identify practical and easily interpretable allocation schedules. Proposition 5.2 below provides these sufficient conditions, for $B=T,N_{t}=1$ for all $t\in[T]$. We defer the proof to Section B.7.
Proposition 5.2.
Suppose there exists $\alpha\in(0,1)$ such that:
  1. 1.
    $\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]>\min\mathopen{}\mathclose{{}\left\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil}\right\}$
  2. 2.
    $\sum_{b}\frac{\mathrm{Var}\mathopen{}\mathclose{{}\left[T_{b}}\right]}{\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]-\min\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\}}\right)^{2}}\leq\frac{1}{2}T^{1-\alpha}$
Then, for any $\delta\geq 3\log(T)e^{-\frac{1}{8}T^{1-\alpha}}$, the process is $\delta$-offset-expiring, and $\underline{X}\geq 1-T^{-\alpha}$.
Proposition 5.2 highlights the two key components that determine the baseline allocation: the distance between the expected perishing time $\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]$ and the expected allocation time $\min\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\}$ (which we colloquially refer to as “room to breathe”), and the variance of the perishing time. Specifically, Condition 1 implies that, if $\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]$ is low, it must be that the item is allocated early on in the horizon (i.e., $\sigma(b)$ is low). This encodes the “race against time” intuition that is typically held around perishing. Condition 2 can be viewed as an upper bound on the cumulative adjusted coefficient of variation (CV) of the perishing process. High-variance perishing times and smaller “room to breathe” push $\alpha$ down, resulting in a lower allocation rate. Hence, to guarantee a high allocation rate, the perishing process needs to satisfy one of two conditions: (1) low variability, or (2) high room to breathe. Having identified these driving factors, we compare the following candidate allocation orders in experiments:
  • Increasing Mean: Increasing order of $\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]$;
  • Decreasing Coefficient of Variation (CV): Decreasing order of $\mathrm{Std}\mathopen{}\mathclose{{}\left[T_{b}}\right]/\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]$. For fixed expected perishing time, this schedule allocates high-variance units earlier on. Conversely, for fixed variance, it allocates items according to the Increasing Mean schedule;
  • Increasing Lower Confidence Bound (LCB): Increasing order of $\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]-1.96\mathrm{Std}\mathopen{}\mathclose{{}\left[T_{b}}\right]$. This ordering allocates items according to the lower bound of the 95% confidence interval of the normal approximation to its perishing time. This lower bound is expected to be small if either the expected perishing time is small or the variance is large.
@@
As in Section 5.1, we draw the demands $N_{t}$ from a truncated normal distribution, $\mathcal{N}(2,0.25)$; we moreover let $T=50$, $B=100$, $\delta=\frac{1}{T}$, and $L_{T}={T}^{-0.35}$. Finally, we consider two sets of perishing distributions:
@@
    $T_{b}=\begin{cases}b+1&\quad b\leq T\\ \text{Uniform}(b+1,T)&\quad b>T\end{cases}$ (18)
@@
It can easily be verified that both instances are $\delta$-offset-expiring, for $\delta=0.05$. Tables 3 and 4 show the performance of our algorithm over these instances.
@@
Order $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathcal{L}^{\textsf{perish}}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Spoilage}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Stockout}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right]$ Increasing Mean $0.10\pm 0.004$ $0.36\pm 0.02$ $0.51\pm 0.02$ $5.78\pm 0.4$ $0.04\pm 0.04$ $6.28\pm 0.4$ Decreasing CV / Increasing LCB $0.0\pm 0$ 0.0 $0.44\pm 0.02$ $0.48\pm 0.02$ $1.24\pm 0.1$ $0.06\pm 0.04$ $1.79\pm 0.1$
Table 3: Performance of Perishing-Guardrail for $L_{T}=T^{-0.35}$ on the distributions given in Eq. 17.
@@
Order $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathcal{L}^{\textsf{perish}}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Spoilage}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Stockout}}\right]$ $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right]$ Increasing Mean / LCB $0.0\pm 0$ 0.0 $0.41\pm 0.02$ $0.46\pm 0.02$ $0.0\pm 0$ 0.0 $0.1\pm 0.05$ $0.56\pm 0.07$ Decreasing CV $0.47\pm 0.0$ $0.51\pm 0.05$ $0.48\pm 0.02$ $48.3\pm 0.08$ $0.01\pm 0.0$ $48.7\pm 0.07$
Table 4: Performance of Perishing-Guardrail for $L_{T}=T^{-0.35}$ on the distributions given in Eq. 18.
@@
For the first instance, the Increasing Mean schedule allocates the first $T$ items uniformly at random, ignoring the fact that, for $b\leq T$, as $b$ increases the item is more likely to perish earlier on in the horizon. The Decreasing CV / Increasing LCB schedules, on the other hand, are identical: they allocate the first $T$ resources in decreasing order of $b$, and allocate the remaining uniformly at random. Notably, the Decreasing CV / Increasing LCB order achieves $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathcal{L}^{\textsf{perish}}}\right]=0$, i.e., $\underline{X}=B/\overline{N}$, as in the no-perishing setting. (Note that $\mathcal{L}^{\textsf{perish}}=0$ implies that this is an optimal ordering.) Since its baseline allocation is higher it results in 78% less spoilage than the Increasing Mean order, and a 71% decrease in inefficiency. However, this order performs slightly worse with respect to counterfactual envy and stockouts: this is again due to the more aggressive allocations.
@@
For the second instance, the Increasing Mean and Increasing LCB schedules are identical: they allocate items lexicographically. The Decreasing CV schedule, on the other hand, allocates the last $T$ items (in increasing order of $b$) before the first $T$ resources, since $\mathrm{Std}\mathopen{}\mathclose{{}\left[T_{b}}\right]=0$ for all $b\leq T$. In this setting, the first schedule is optimal with respect to $\sigma$-induced loss, with $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathcal{L}^{\textsf{perish}}}\right]=0$. This more aggressive allocation results in a 10% stockout rate (versus 1% for the Decreasing CV schedule), but outperforms the Decreasing CV order across all other metrics. This is intuitive as the number of errors in this latter, clearly bad order results in $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathcal{L}^{\textsf{perish}}}\right]=0.47$, approximately $50\%$ of the baseline allocation $B/\overline{N}$. The algorithm then incurs both high inefficiency and spoilage.

### v2

@@ 5.2 Geometric Perishing: Synthetic Experiments
Figure 6: Algorithm comparison across $\mathbb{E}\left[\Delta_{\text{\it EF}}\right],\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right],\mathbb{E}\left[\textsc{Envy}\right]$, and $\mathbb{E}\left[\textsc{Spoilage}\right]$, for $\alpha\in\{0.1,0.2,0.25,0.3\}$.
We identify three regimes:
  • High Perishing ($\alpha=0.1$, Figure 6(a)): While unfairness (as measured by $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ and $\mathbb{E}\left[\textsc{Envy}\right]$) is decreasing in $T$ under Perishing-Guardrail and Static-$\underline{X}$, Vanilla-Guardrail and Static-$\frac{B}{\overline{N}}$ perform remarkably poorly along these two metrics. This is due to the fact that these latter algorithms fail to account for the unavoidable perishing loss, resulting in an extremely high stockout probability, as illustrated in Table 1, for $T=150$. In contrast, the two perishing-aware algorithms rarely run out of resources. This underscores the importance of modifying the baseline guardrail $\underline{X}$, which was specifically constructed to avoid stockouts due to unavoidable perishing.
    Comparing Static-$\underline{X}$ to Perishing-Guardrail, our results also demonstrate that, in this high-perishing regime, the strategy of cautiously over-allocating by $L_{T}$ comes at a significant reduction in inefficiency $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$, at close to no increase in counterfactual envy $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$.
  • Medium Perishing ($\alpha=0.2$, Figure 6(b)): Though we observe similar trends as when $\alpha=0.1$, all algorithms perform better across the board. Perishing-Guardrail exhibits slightly higher efficiency and lower spoilage than its perishing-agnostic counterpart in this regime, since it satisfies the threshold condition more frequently, as described above. Still, in Table 1 we see that the perishing-agnostic algorithms run out of resources in over 50% of replications.
  • Low Perishing ($\alpha=0.3$, Figure 6(c)): For this smaller perishing rate, Vanilla-Guardrail stocks out significantly less frequently. Putting this together with the fact that $B/\overline{N}>\underline{X}$, this explains the fact that it has lower counterfactual envy than Perishing-Guardrail. However, along all other metrics Perishing-Guardrail improves upon Vanilla-Guardrail. The improvements in efficiency and spoilage are due to the same effects as described above; moreover, our algorithm improves upon Vanilla-Guardrail on $\mathbb{E}\left[\textsc{Envy}\right]$ since it never stocks out.
$\alpha$ $\frac{B}{\overline{N}} Static$-$\underline{X}$ Vanilla-Guardrail Perishing-Guardrail
$0.1$ $99\pm 0.02 0.0$$00\pm 0.0$ $0\pm 0.0 0.11±0$$1\pm 0.06$
$0.2$ $63\pm 0.1 0.0$$00\pm 0.0$ $68\pm 0.09 0.0$$03\pm 0.03$
$0.25$ $17\pm 0.07 0.0$$00\pm 0.0$ $22\pm 0.08 0.0$$00\pm 0.0$
$0.3$ $3\pm 0.04 0.00±0$$\pm 0.0 0.0$$6\pm 0.05 0.0$$00\pm 0.0$
@@
Overall, our results highlight the robustness of Perishing-Guardrail to perishability, as it is able to achieve similar if not improved performance compared to Vanilla-Guardrail in settings where there is limited perishing, with vastly superior performance in high-perishing settings.
Algorithm $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ $\mathbb{E}\left[\textsc{Envy}\right]$ $\mathbb{E}\left[\textsc{Spoilage}\right]$ $\mathbb{E}\left[\textsc{Stockout}\right]$ $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$
@@
Table 2: Performance of the different algorithms (for $L_{T}=T^{-0.35}$) on the “ginger” dataset in Keskin et al. (2022). The second number in each cell corresponds to 95% confidence intervals.
5.3 Geometric Perishing: Real-World Instance
We next investigate the performance of our algorithm using the real-life dataset provided by Keskin et al. (2022), which tracks daily demand, replenishments, and perishing of ginger at a leading supermarket chain in China over $T=365$ days. We assume each unit’s perishing time is geometrically distributed with parameter $p$. Fitting this geometric perishing model to the data, we obtain an estimate of $p=0.00224$. We similarly fit a truncated normal distribution to the demand data, obtaining $\mathcal{N}(3.2,1.85^{2})$. Finally, we set the total budget to $B=365\cdot 3.2=1168$, corresponding to the total expected demand over the horizon. (See Section 9 for additional details on the data calibration.) For these inputs, $B/\overline{N}=0.89$ and $\underline{X}=0.46$. Under these parameters, the offset-expiry condition is only satisfied 65.2% of the time. Given this aggressive perishing, perishing-agnostic algorithms are expected to perform particularly poorly.
@@ 5.3 Geometric Perishing: Real-World Instance
  • As conjectured, Static-$\frac{B}{\overline{N}}$ and Vanilla-Guardrail stock out on 100% of replications since they fail to account for endogenous perishing. The high stockout probabilities of Static-$\frac{B}{\overline{N}}$ and Vanilla-Guardrail lead to high unfairness (vis-à-vis hindsight and counterfactual envy), since later arrivals receive allocations of zero. In contrast, Static-$\underline{X}$ never stocks out. Perishing-Guardrail achieves a higher stockout rate of 40%, due to its more aggressive allocations. Still, our algorithm’s counterfactual envy and hindsight envy are over 30% and 70% lower, respectively, than that of Vanilla-Guardrail.
  • Perishing-Guardrail allocates approximately 10% fewer goods than Vanilla-Guardrail. It is notable, however, that it is more efficient than Static-$\frac{B}{\overline{N}}$; this highlights that naively allocating more aggressively need not always generate efficiency gains.
Overall, we see that even when offset-expiry fails to hold over a significant fraction of sample paths, for small losses in efficiency our algorithm makes major gains in fairness relative to perishing-agnostic algorithms. This highlights the robustness of our algorithm to aggressive perishing.
@@
Theorem 4.1 established that the theoretical performance of Perishing-Guardrail improves for orderings $\sigma$ under which $\mathcal{L}^{\textsf{perish}}$ is lower. In practice, however, $\mathcal{L}^{\textsf{perish}}$ has an extremely subtle impact on our algorithm’s efficiency. Namely, while lower $\mathcal{L}^{\textsf{perish}}$ implies a higher $\underline{X}$, this also results in a lower value of $\overline{P}_{t}$ in each round $t$. The net effect on whether the threshold condition is satisfied in each period, then, remains unclear. In this section we study the impact of reasonable heuristic orderings on the practical performance of our algorithm. Motivated by the insights of Proposition 4.19, we compare the following orderings:
  • Increasing Mean ($\sigma^{\textsf{Mean}}$): Increasing order of $\mathbb{E}\left[T_{b}\right]$;
  • Decreasing Coefficient of Variation (CV) ($\sigma^{\textsf{CV}}$): Decreasing order of $\mathrm{Std}\left[T_{b}\right]/\mathbb{E}\left[T_{b}\right]$. For fixed expected perishing time, this schedule allocates high-variance units earlier on. Conversely, for fixed variance, it allocates items according to the Increasing Mean schedule;
  • Increasing Lower Confidence Bound (LCB) ($\sigma^{\textsf{LCB}}$): Increasing order of $\mathbb{E}\left[T_{b}\right]-1.96\mathrm{Std}\left[T_{b}\right]$. This ordering allocates items according to the lower bound of the 95% confidence interval of the normal approximation to its perishing time. This lower bound is expected to be small if either the expected perishing time is small or the variance is large.
@@
As in Section 5.2, we draw the demands $N_{t}$ from a truncated normal distribution, $\mathcal{N}(2,0.25)$; we moreover let $T=50$, $B=100$, and $L_{T}={T}^{-0.35}$. Finally, we consider two sets of perishing distributions:
@@
    $T_{b}=\begin{cases}b+1&\quad b\leq T\\ \text{Uniform}(T,b+1)&\quad b>T\end{cases}$ (24)
@@
Tables 3 and 4 show the performance of our algorithm over these instances, for all three allocation schedules.
@@
Order $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]$ $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ $\mathbb{E}\left[\textsc{Envy}\right]$ $\mathbb{E}\left[\textsc{Spoilage}\right]$ $\mathbb{E}\left[\textsc{Stockout}\right]$ $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$
$\sigma^{\textsf{Mean}}$ $0.10\pm 0.004$ $0.36\pm 0.02$ $0.51\pm 0.02$ $5.78\pm 0.4$ $0.04\pm 0.04$ $6.28\pm 0.4$
$\sigma^{\textsf{CV}}$ / $\sigma^{\textsf{LCB}}$ $0.0\pm 0.0$ $0.44\pm 0.02$ $0.48\pm 0.02$ $1.24\pm 0.1$ $0.06\pm 0.04$ $1.79\pm 0.1$
Table 3: Performance of Perishing-Guardrail for $L_{T}=T^{-0.35}$ on Instance 1 (Eq. 23).
@@
Order $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]$ $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ $\mathbb{E}\left[\textsc{Envy}\right]$ $\mathbb{E}\left[\textsc{Spoilage}\right]$ $\mathbb{E}\left[\textsc{Stockout}\right]$ $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$
$\sigma^{\textsf{Mean}}$ / $\sigma^{\textsf{LCB}}$ $0.0\pm 0.0$ $0.41\pm 0.02$ $46\pm 0.02 0.0$$0\pm 0.0$ $0.1\pm 0.05$ $0.56\pm 0.07$
$\sigma^{\textsf{CV}}$ $0.47\pm 0.0$ $0.51\pm 0.05$ $0.48\pm 0.02$ $48.3\pm 0.08$ $0.01\pm 0.0$ $48.7\pm 0.07$
Table 4: Performance of Perishing-Guardrail for $L_{T}=T^{-0.35}$ on Instance 2 (Eq. 24).
@@
For the first instance, the Increasing Mean schedule allocates the first $T$ items uniformly at random, ignoring the fact that, for $b\leq T$, as $b$ increases the item is more likely to perish earlier on in the horizon. The Decreasing CV / Increasing LCB schedules, on the other hand, are identical: they allocate the first $T$ resources in decreasing order of $b$, and allocate the remaining uniformly at random. Notably, the Decreasing CV / Increasing LCB order achieves $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]=0$, i.e., $\underline{X}=B/\overline{N}$, as in the no-perishing setting. (Note that $\mathcal{L}^{\textsf{perish}}=0$ implies that this is an optimal ordering.) Since its baseline allocation is higher it results in 78% less spoilage than the Increasing Mean order, and a 71% decrease in inefficiency. However, this order performs slightly worse with respect to counterfactual envy and stockouts: this is again due to the more aggressive allocations.
@@
For the second instance, the Increasing Mean and Increasing LCB schedules are identical: they allocate items lexicographically. The Decreasing CV schedule, on the other hand, allocates the last $T$ items (in increasing order of $b$) before the first $T$ resources, since $\mathrm{Std}\left[T_{b}\right]=0$ for all $b\leq T$. In this setting, the first schedule is optimal with respect to perishing-induced loss, with $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]=0$. This more aggressive allocation results in a 10% stockout rate (versus 1% for the Decreasing CV schedule), but outperforms the Decreasing CV order across all other metrics. This is intuitive as the number of errors in this latter, clearly bad order results in $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]=0.47$, approximately $50\%$ of the baseline allocation $B/\overline{N}$. The algorithm then incurs both high inefficiency and spoilage.